how long will it take for an investment to double if it earns 6.25% interest per year compounded quarterly?
Answer by mananth(16949)

Principal P = 100
Amount= 200
n= number of periods
compounded (q) 4
Rate = 6.25
n = ln(FV / PV) / ln(1 + r/q)

n=( ln ( 200 / 100 )/ln( 1 + 0.02 )
n=( 0.69 )/ ln 1.02
n=( 0.69 / 0.02 )
n= 44.71 quarters
divide by 4
11.18 years